a rollover ira of $18,750 was invested in two mutual funds one earning 12% interest and the other earning 10% after 1 year, the combined interest income is 2,113 how much was invested at each rate
Answer by mananth(16949)

Investment I 12.00% per annum
Investment II 10.00% per annum

x + y 18750 ------------------------1
12.00% x + 10.00% y = 2113
Multiply by 100
12 x + 10 y = 211300 --------2

Multiply (1) by -10
we get

-10 x -10 y = -187500.00

Add this to (2)

2 x 0 y = 23800.00

divide by 2

y = 11900 investment at 10.00%
Balance 6850 investment at 12.00%